OpenStudy (anonymous):

Find the derivative of the function. h(x) = 7x^2 − x

OpenStudy (anonymous):

\[Thats 7x ^{2-x}\]

OpenStudy (anonymous):

14x-1

OpenStudy (anonymous):

oh nvm lol use chain rule

OpenStudy (anonymous):

Is it ln(7)(7)\[x^{2-x}\]

OpenStudy (anonymous):

include x^2-x after 7

OpenStudy (anonymous):

i believe its -x^(1-x)

OpenStudy (anonymous):

@Spidermonkeyy im pretty sure it should look similar to \[\ln(7)(7)^{x ^{2}-x}\]

OpenStudy (anonymous):

because thats what all the examples look like

OpenStudy (anonymous):

i used chain rule but im not sure about lns in there sorry

OpenStudy (shamil98):

not h(x) = 7x^2 -x its \[7x^{2-x}\]

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

\[h(x) = 7x^{2 -x} \] taking log both sides we find: \[\log h(x) = \log(7x^{2 -x}) \rightarrow \log h(x) = \log7 + \log x^{2 -x}\] \[\rightarrow \log h(x) = \log7 +(2-x) \log x\] Differentiating with respect to x we find \[\frac{1}{h(x)} h'(x) = 0 + (2-x) \frac{1}{x} + \log x (0-1)\] \[\rightarrow \frac{1}{h(x)} h'(x) = \frac{(2-x)}{x} + \log x (-1)\] \[\rightarrow h'(x) = h(x) (\frac{(2-x)}{x} - \log x)\] \[\rightarrow h'(x) = 7x^{2-x} (\frac{(2-x)}{x} - \log x)\] \[\rightarrow h'(x) = 7x^{2-x} (\frac{(2-x-x \log x)}{x} )\] \[\rightarrow h'(x) = 7x^{2-1-x} (2-x-x \log x)\] \[\rightarrow h'(x) = 7x^{1-x} (2-x-x \log x)\] @arshia93

OpenStudy (anonymous):

@arshia93 Your question has been solved.....

OpenStudy (anonymous):

@dpasingh bravo! Can you please answer a different question I have if you cant find it ill bump it in 7 mins